OpenStudy (nilankshi):

obtain all the zeroes of 3x^4+6x^2-2x^2-10x-5,if two zeroes are root5/3 and -root 5/3

OpenStudy (anonymous):

I assume you mean: 3x^4+6x^3-2x^2-10x-5

OpenStudy (lgbasallote):

the roots aree \(\large \sqrt{\frac{5}{3}}\) and \(\large -\sqrt{\frac{5}{3}}\) right? or is the root applied on 5 only?

OpenStudy (anonymous):

is there a square root sign at all? I understood that 5/3 and -5/3 are the roots

OpenStudy (apoorvk):

divide the whole system by its two factors that are given (or try synthetic division) After two whole divisions by each factor, you 'll be left with a quadratic as the quotient, factorising which you'll get you the other two roots.

OpenStudy (lgbasallote):

he/she said root 5/3 and root -5/3..she already said zeroes so i assume root->radical

OpenStudy (nilankshi):

yes that is right..

OpenStudy (nilankshi):

root is on both..

OpenStudy (apoorvk):

\[\mathbb{yes}\]

OpenStudy (anonymous):

and is it an x^3 in the equation?

OpenStudy (anonymous):

(6x^2) should be 6x^3

OpenStudy (nilankshi):

yes it is in equion

OpenStudy (lgbasallote):

\(\large x = \sqrt{\frac{5}{3}}\) \(\large x^2 = \frac{5}{3}\) \(\large 3x^2 = 5\) \(\large 3x^2 - 5 = 0\) as for the other zero... \(\large x = -\sqrt{\frac{5}{3}}\) \(\large x^2 = \frac{5}{3}\) \(\large 3x^2 = 5\) \(\large 3x^2 - 5 = 0\) so two of the factors would be (3x^2 - 5)^2

OpenStudy (lgbasallote):

something doesnt feel right...

OpenStudy (nilankshi):

yaa..

OpenStudy (lgbasallote):

hmm idk how to do this sorry... @apoorvk may know though

OpenStudy (nilankshi):

its ok..

OpenStudy (apoorvk):

Sorry, Trance. :) and Nilakshi. I 'll work on the solution now

OpenStudy (trancenova):

Thanks :) @Rohangrr if you want to talk about it, please make a new question.

OpenStudy (phi):

I would multiply out \[ (x-\sqrt{5/3})(x+\sqrt{5/3}) = x^2-\frac{5}{3}\] we can write this as \(3x^2-5\) (why? see below) divide this into the original equation to get \(x^2+2x+1 \) now factor this to get the remaining roots. * we are looking for \((x^2-\frac{5}{3}) (x^2+2x+1) =0\) if we multiply both sides by 3 we get \( 3(x^2-\frac{5}{3})(x^2+2x+1) =0\) \( (3x^2-5) (x^2+2x+1)=0\)

OpenStudy (apoorvk):

Hmm, I got so bored, Phi already posted it. :/ Good job, Phi